# How Many Van Goghs Does It Take to Van Gogh? Finding the Imitation Threshold

### v1

@@ 4 Experimental Setup
Table 1: Domains, datasets, pretraining data,
and models we experiment with.
Domain Dataset Pretraining Data Model Human Faces [Uncaptioned image] Celebrities, Politicians LAION2B-en SD1.1, SD1.5 Human Faces [Uncaptioned image] Celebrities, Politicians LAION5B SD2.1 Art Style [Uncaptioned image] Classical, Modern LAION2B-en SD1.1, SD1.5 Art Style [Uncaptioned image] Classical, Modern LAION5B SD2.1
Table 2: Five prompts used to generate images of human faces and art styles.
@@
Samara Joy    Jacquetta Wheeler    James Garner     Stephen King     Johnny Depp
@@
Text-to-image Models and Training Data. We use Stable Diffusion (SD) as the text-to-image models Rombach et al. (2022). We use them because both the models and their training datasets are open-sourced. Specifically, we use SD1.1 and SD1.5 that were trained on LAION2B-en, a 2.3 billion image-caption pairs dataset, filtered to contain only English captions. In addition, we use SD2.1 that was trained on LAION-5B, a 5.85 billion image-text pairs dataset, which includes LAION2B-en, and other image-caption pairs from other languages Schuhmann et al. (2022).
@@
Domains and Concepts. We experiment with two domains – art styles [Uncaptioned image] and human faces [Uncaptioned image] that are highly important for privacy and copyright considerations of text-to-image models. Figures 1 and 2 show examples of real and generated images of art styles and human faces. We collect two sets of concepts for each domain. For art styles we collect classical and modern art styles, and for human faces, we collect celebrities and politicians. These sets are independent (i.e., have no common concept) and are therefore useful to test the robustness of the thresholds (§6). Each set has 400 concepts that cover a wide frequency range in the pretraining datasets. Appendix L provides details of the sources used to collect the concepts and the sampling procedure. Table 2 summarizes the domains, sets used for each domain, the models and their pretraining datasets we experiment with.
@@
Image Generation. We generate images for each domain by prompting models with five prompts (Table 2). We design domain-specific prompts that encourage the concepts to occupy most of the generated image, which simplifies the imitation score measurement. We also ensure that these prompts are distinct from the captions used in the pretraining dataset to minimize reproduction of training images. (Somepalli et al. (2023b) proposed this strategy to mitigate reproduction of training images.) Out of the five prompts used to generate images, only one of them was found in the LAION2B-en dataset caption 3 times out of the total 2.3B captions. We generate 200 images per concept using different random seeds for each prompt, a total of 1,000 images per concept.
5 Proposed Methodology: MIMETIC2
We illustrate our proposed methodology in Figure 3. At a high level, for a specific domain, MIMETIC2 estimates the frequency of each concept in the pretraining data (Section 5.1) and estimates the model’s ability to imitate it (Section 5.2). We then sort the concepts based on their frequencies, and find the imitation threshold using a change detection algorithm (LABEL:{subsec:detecting-threshold}).
5.1 Concept Frequency
Challenges.
Figure 4: LAION images whose caption mentions ‘Mary Lee Pfeiffer’, the mother of Tom Cruise. She is not always present in the images (e.g., the rightmost image).
Determining a concept’s frequency in a multimodal dataset can be achieved by employing a high-quality classifier for that concept over every image and counting the number of detected images. However, given the scale of modern datasets with billions of images, this approach is expensive and time consuming. Instead, we make a simplifying assumption that a concept is present only if the image’s caption mentions it. We further discuss this assumption and provide supporting evidence for its accuracy in Appendix C. In addition, concepts often do not appear in the corresponding images, even when they are mentioned in the captions. For instance, Figure 4 showcases images whose captions contain “Mary Lee Pfeiffer”, but such images do not always include her. On average, we find that concepts occur only in 60% of the images whose caption mentions the concept.
Estimating Concept Frequency We start by retrieving all images whose caption mentions the concept of interest and then further filter out the images that do not contain the concept, as detected by a classifier. We retrieve these images using WIMBD (Elazar et al., 2024), a search tool based on a reverse index that efficiently finds documents (captions) from LAION containing the search query (concept). In addition, for each concept, we construct a set of high quality reference images. For example, a set of images with only the face of a single person (e.g., Brad Pitt). We collect these images automatically using a search engine, followed by a manual verification to vet the images (see Appendix D for details). These images are used as gold reference for automatic detection of these concepts in the images from the pretraining datasets (we collect up to ten reference images per concept).
To classify whether a candidate image from the pretraining data contains the concept of interest, we embed the candidate image and the concept’s reference images using an image encoder and measure the similarity between the embeddings. We use a face embedding model (Deng et al., 2022) for faces and an art style embedding model (Somepalli et al., 2024) for art style. If the similarity between a candidate image and any of the reference images is above some threshold, we consider that candidate image to contain the concept. This threshold is established by measuring the similarity between images of the same concepts and images of different concepts which maximizes the true positives, and minimizes false positives. We provide additional details on how we find the exact thresholds for both art styles and human faces in Appendices F and E.
Finally, we employ the classifier on all candidate images corresponding to a concept (i.e., images whose caption mentions the concept), and take those that are classified as positive. For each concept, we randomly retrieve up to 100K candidate images. We use the ratio of positive predictions to the total number of retrieved candidate images, and multiply it by the total caption count of a concept in the dataset and use that as the concept’s frequency estimate. For concepts with less than 100K candidate images, we simply use all the images that are positively classified. Note that several URLs in the LAION datasets are dead, a common phenomenon for URL based datasets (“link rot” Carlini et al. (2024); Lakic et al. (2023)). On average, we successfully retrieved 74% of the candidate images.
5.2 Computing Imitation Score
Challenges. Computing the imitation score entails determining how similar a concept is in a generated image compared to its source images from the training data. Several approaches were proposed to accomplish this task, such as FID and CLIPScore Salimans et al. (2016); Heusel et al. (2017); Sajjadi et al. (2018); Hessel et al. (2021); Podell et al. (2024). To measure similarity, these approaches compute the similarity between the distributions of the embeddings of the generated and training images of a concept. The embeddings are obtained using image embedders like Inception model in case of FID and CLIP in case of CLIPScore. These image embedders often perform reasonably well in measuring similarity between images of common objects which constitutes most of their training data. However, they cannot reliably measure the similarity between two very similar concepts like the faces of two individuals or art style of two artists (Somepalli et al., 2024; Hessel et al., 2021; Ahmadi & Agrawal, 2024; Jayasumana et al., 2024). Therefore, MIMETIC2 uses domain specific image embedders to measure similarity between two concepts from that domain. It uses a face embedding model (Deng et al., 2022) for measuring face similarity and an art style embedding model Somepalli et al. (2024) for measuring art style similarity. Even the specific choice of these models is crucial. For instance, in early experiments we used Facenet (Schroff et al., 2015), and observe it struggles to distinguish between individuals of certain demographics, causing drastic differences in the imitation scores between demographics. We provide more details on these early experiments in Appendix K, and show that our final choice of embedding models work well on different demographics.
Estimating Imitation Score To measure the imitation score we embed the generated images and filtered training images of a concept (obtained in §5.1) using the domain specific image embedder, and report the imitation score as the average imitation between all generated and training images. For measuring face imitation, we use InsightFace, a face embedding model (Deng et al., 2020) that extracts the individual’s face from an image and generates an embedding for it. For measuring art style imitation, we use CSD, an art style embedding model (Somepalli et al., 2024) that generates an embedding for an image of an art work. We obtain the embeddings of the generated and training images for each concept, and measure imitation by computing the cosine similarity between them.
To ensure that the automatic measure of similarity correlates with human perception, we also conduct experiments with human subjects and measure the correlation between the similarities obtained automatically and in the human subject experiments. We find high correlation between the automatically obtained scores and human perception of similarity for both domains (§6).
5.3 Detecting the Imitation Threshold
After computing the frequencies and the imitation scores for each concept, we sort them in an ascending order of their image frequencies. This generates a sequence of points, each of which is a pair of image frequency and the imitation score of a concept. We apply a standard change detection algorithm, PELT (Killick et al., 2012), to find the image frequency where the imitation score significantly changes. Change detection is a classic statistical approach for which the objective is to find the points where the mean value of a sequence changes significantly. It is used for detecting changes such as shifts in stock market trends or network traffic anomalies. Several algorithms were proposed for change detection (van den Burg & Williams, 2022). We choose PELT because of its linear time complexity in computing the change point. We choose the first change point as the imitation threshold (see Appendix H for details about all change points). The application of change detection assumes that increasing the image counts beyond a certain threshold leads to a large jump in the imitation scores, and we find this assumption to be accurate in our experimental results.
@@ 6 Results: The Imitation Threshold
Table 3: Imitation Thresholds for human face and art style imitation for the different text-to-image models and pretraining datasets we experiment with.
Human Faces [Uncaptioned image] Art Style [Uncaptioned image] Pretraining Dataset Model Celebrities Politicians Classical Modern LAION2B-en SD1.1 364 234 112 198 SD1.5 364 234 112 198 LAION-5B SD2.1 527 369 185 241
@@
We apply MIMETIC2 to estimate the imitation threshold for each model-data pair, and present the results in Table 3. The imitation thresholds for SD1.1 on celebrities and politicians are 364 and 234 respectively. And the imitation thresholds for classical and modern art styles are 112 and 198 respectively. Interestingly, SD1.1 and SD1.5 have the same thresholds for all the four sets across both the domains. Notably, both SD1.1 and SD1.5 are trained on LAION2B-en. The imitation thresholds for SD2.1, which is trained on the larger LAION-5B dataset is higher than the thresholds for SD1.1 and SD1.5. The imitation threshold for SD2.1 on celebrities and politicians are 527 and 369 respectively, and on classical and modern art styles are 185 and 241 respectively. We hypothesize that the difference in performance of SD2.1 and SD1.5 is due to the difference in their text encoders (O’Connor, 2022) (note: differences in performance between SD1.5 and SD2.1 were also reported by several users on online forums). To test this hypothesis, we compute the imitation thresholds for politicians for all SD models in series 1: SD1.1, SD1.2, SD1.3, SD1.4, and SD1.5. We find that the imitation thresholds for all these models are almost the same. We provide the imitation thresholds for all series 1 models in Appendix G.
Note that celebrities have a higher imitation threshold than politicians. We hypothesize this happens due to inherent differences in the data distribution between these two sets, which makes it harder to learn the concept of celebrities than politicians. To test this hypothesis, we compute the average number of images that has one person for people with less than 1,000 images in the pretraining dataset. We find that politicians have about twice the number of single person images compared to celebrities. As such, images that have a single person (the concept of interest) increase the ability of the model to learn from them, thus lowering the imitation threshold for politicians.
We also present the plots of the imitation scores as a function of the image frequencies of the concepts in the four sets we experiment with. Figures 5(a) and LABEL:{fig:artists-group1-plot-similarity-SD1.1} show the imitation graphs of celebrities and classical art styles, respectively for SD1.1. The x-axis is the frequency and the y-axis is the imitation score (averaged over the five image generation prompts) for each concept, sorted in increasing order of frequency. We showcase the imitation graphs for all other models and sets in Appendix I, which follow similar trends.
(a) Human Face Imitation. The imitation threshold is detected at 364 faces.
(b) Art Style Imitation. The imitation threshold is detected at 112 images.
Figure 5: Human Face and Art Style imitation graphs for SD1.1 using the Celebrities and Classical art style sets. The x-axis represents the sorted image frequencies in the training dataset, and the y-axis represents the imitation of the training images in the generated images, for each concept. Concepts with zero image frequencies are shaded with light gray. We show the mean imitation score and its variance over the five image generation prompts. The red vertical line indicates the imitation threshold found by the change detection algorithm, and the horizontal green line represents the average imitation scores before and after the threshold.
@@
In Figure 5(a), we observe that the imitation scores for individuals with low image frequencies are close to 0 (left side), and increase as the image frequencies increase (towards the right side). The highest similarity is 0.5 and it is for individuals in the rightmost region of the plot. We observe a low variance in the imitation scores across prompts, and also note that this variance does not depend on the image frequencies ($0.0003\pm 0.0005$) – indicating that the performance of the face embedding model does not depend on the popularity of the individual.
Similarly, in Figure 5(b), we observe that imitation scores for art styles with low image frequencies are low (close to 0.2 on the left side), and increase as the image frequencies increase (towards the right side). The highest similarity is 0.76 and it is for the artists in the rightmost region of the plot. We also observe a low variance across the generation prompts, and the variance does not depend on the popularity of the artist ($0.003\pm 0.003$).
@@
Human Perception Evaluation. To determine if the automatic measure of similarity between the generated and training images correlate with human perception, we conduct experiments with human subjects. We asked the participants to rate generated images on the Likert scale Likert (1932) of 1-5 based on their similarity to real images of celebrities (the same ones used for measuring the imitation score). To avoid any bias, the participants were not informed of the research objective of this work.
@@
For human face imitation, we conduct this study with 15 participants who were asked to rate 100 (randomly selected) generated images for a set of 40 celebrities. To determine the accuracy of the imitation threshold estimated by MIMETIC2, we select the celebrities such that half of them have image frequencies below the threshold and the other half above it. We measure the Spearman’s correlation (Zar, 2005) between the imitation scores computed by the model and the ratings provided by the participants. Due to the variance in perception, we normalize the ratings for each participants. The Spearman correlation between the similarity scores provided by participants and the imitation scores is 0.85, signifying a high quality imitation estimator. We also measure the agreement between the imitation threshold that MIMETIC2 estimates and the threshold that humans perceive. For this purpose, we convert the human ratings to binary values and treat it as the ground truth (any rating of 3 or more is treated as 1 and less than 3 is treated as 0). As for the MIMETIC2’ predictions, we construct another set of the same size that has a zero for a celebrity whose concept frequency is lower than the imitation threshold, and 1 otherwise. To measure the agreement, we compute the element-wise dot product between these two sets. We find the agreement to be 82.5%, signifying a high degree of agreement for MIMETIC2’s automatically computed threshold.
@@
Results Discussion. Overall, we observe that the imitation thresholds are similar across the different image generation models and pretraining datasets, but are domain dependent. They also show little variance across various image generation prompts. Most importantly, the thresholds computed by MIMETIC2 have a high degree of agreement with human subjects.
@@
We also note the presence of several outliers in both plots, which can be categorized into two types: (1) concepts whose image frequencies are lower than the imitation threshold, but their imitation scores are considerably high; and (2) concepts whose image frequencies are higher than the imitation threshold, but their imitation scores are low. From a privacy perspective, the first kind of outliers are more concerning than the second ones, since the imitation threshold should act as guarantors of privacy. It would be fine if a concept with a concept frequency higher than the threshold is not imitated by the model (false positive), but it would be a privacy violation if a model can imitate a concept with frequency lower than the threshold (false negative). Therefore, to minimize false negatives, it is preferable to underestimate the imitation threshold. Upon further analysis, we find that the actual concept frequencies of all the false negative outliers is much higher than what MIMETIC2 counts, primarily due to aliases of names, thereby alleviating the privacy violation concerns (see §7).
@@
We also note that the range of the imitation scores of different domains have different y-axis scales. This is due to the difference in embedding models used in both cases. The face embedding model can distinguish between two faces much better than the art style model can distinguish between two styles (see Appendices E and F), and therefore the scores for the concepts on the left side of the imitation threshold is around 0 for face imitation and 0.2 for style imitation. The face embedding model also gives lower score to the faces of the same person, compared to the style embedding model’s score for images of the same art styles, and therefore the highest scores for face imitation is 0.5, whereas it is 0.76 for art style imitation. However, the absolute values on the y-axis do not matter for estimating the imitation threshold as long as the trend is similar, which is the case for both domains.
Evaluation of the Imitation Thresholds. Evaluating the accuracy of the imitation thresholds MIMETIC2 finds is a very expensive task, it requires conducting the optimal experiment described in Section 3 (training $\mathcal{O}(\log m)$ models with different number of images of each concept). Therefore we choose an approximate method to test MIMETIC2’s accuracy. Concretely, we calibrate MIMETIC2 using one of the sets in each domain (Celebrities for Human Faces [Uncaptioned image] and Classical art styles for Art Style [Uncaptioned image]) and test the calibrated approach on the other set (Politicians for Human Faces [Uncaptioned image] and Modern art styles for Art Style [Uncaptioned image]). In essence, the sets that we use to calibrate MIMETIC2 on, acts as the “train datasets” and the other sets acts as the “test dataset”. Note that these sets are mutually exclusive. Table 3 shows that the imitation thresholds MIMETIC2 finds for “train” and “test” sets in each domain are very close (the difference across the sets for both human faces and art styles is less than 5e-10% of the pretraining dataset size), indicating that the imitation thresholds MIMETIC2 finds are accurate. Note that this is another evidence for MIMETIC2’s accuracy in addition to the evidence already provided by the human subject experiments.

### v2

@@ 5 Experimental Setup
Table 1: Domains, datasets, pretraining data, and models we use.
Domain Dataset Pretraining Data Model Human Faces [Uncaptioned image] Celebrities, Politicians LAION400M LDM Human Faces [Uncaptioned image] Celebrities, Politicians LAION2B-en SD1.1, SD1.5 Human Faces [Uncaptioned image] Celebrities, Politicians LAION5B SD2.1 Art Style [Uncaptioned image] Classical, Modern LAION400M LDM Art Style [Uncaptioned image] Classical, Modern LAION2B-en SD1.1, SD1.5 Art Style [Uncaptioned image] Classical, Modern LAION5B SD2.1
Table 2: Prompts used to generate images of human faces and art styles.
@@
Samara Joy         Jillie Mack         James Garner         Stephen King         Johnny Depp
@@
Table 3: The mean and std. error of the imitation thresholds for the models and domains we experiment with.
Human Faces [Uncaptioned image] Art Style [Uncaptioned image] Pretraining Model Celebrities Politicians Classical Modern LAION-400M LDM 626 $\pm$ 3 449 $\pm$ 2 416 $\pm$ 3 412 $\pm$ 2 LAION2B-en SD1.1 399 $\pm$ 3 284 $\pm$ 3 304 $\pm$ 2 208 $\pm$ 1 SD1.5 371 $\pm$ 1 302 $\pm$ 4 302 $\pm$ 2 212 $\pm$ 1 LAION-5B SD2.1 617 $\pm$ 5 385 $\pm$ 5 330 $\pm$ 4 292 $\pm$ 3
@@
Text-to-image Models and Training Data We use Stable Diffusion (SD) as the text-to-image models (stable-diffusion-model-paper). We use them because both the models and their training datasets are open-sourced. Specifically, we use SD1.1 and SD1.5 that were trained on LAION2B-en, a 2.3 billion image-caption pairs dataset, filtered to contain only English captions and we use SD2.1 that was trained on LAION-5B, a 5.85 billion image-text pairs dataset that includes captions in any language (laion-5b-paper). Finally, we also use a latent diffusion model (LDM) trained on LAION-400M (schuhmann2021-laion400m), a 400M image-text pairs dataset.
@@
Domains and Concepts We experiment with two domains – art styles [Uncaptioned image] and human faces [Uncaptioned image] that are highly important for privacy and copyright considerations of text-to-image models. Figures 1 and 4 show examples of real and generated images of art styles and human faces. We collect two sets of concepts for each domain. For art styles we collect classical and modern art styles and for human faces, we collect celebrities and politicians. It is important to note that each artist’s style and each human face is considered as a separate concept. Although in the real world art style of some artists might be similar, the classifier models we use to measure imitation is trained well to distinguish between art styles from different artists (somepalli2024-style-similarity). These sets we experiment with are independent (i.e., have no common concept) and are therefore useful to test the robustness of the thresholds (§6). Each set has 400 concepts that cover a wide frequency range in the pretraining datasets. Appendix˜N provides details of the sources used to collect the concepts and the sampling procedure. Table˜2 summarizes the domains, sets used for each domain, the models and their pretraining datasets we experiment with.
@@
Image Generation We generate images for each domain by prompting models with five prompts (Table˜2). We design domain-specific prompts that encourage the desired concept to occupy a large part of the generated image, which simplifies the imitation score measurement. We also ensure that these prompts are distinct from the captions used in the pretraining dataset to minimize direct reproduction of training images (as noted by (somepalli2023understanding-and-mitigating)). We generate 200 images per concept using different random seeds for each prompt, a total of 1,000 images per concept.
@@ 6 Results: The Imitation Threshold
We use MIMETIC2 to estimate the imitation threshold for each model-data pair, and present the results in Table˜3. The imitation thresholds for SD1.1 on celebrities and politicians are 399 $\pm$ 85 and 284 $\pm$ 87 respectively, and 304 $\pm$ 56 and 208 $\pm$ 26 for classical and modern art styles, respectively. Interestingly, SD1.1 and SD1.5 have almost identical thresholds across the four datasets. Notably, both SD1.1 and SD1.5 are trained on LAION2B-en. The imitation thresholds for SD2.1, which is trained on LAION-5B (a superset of LAION-2B) are higher than the thresholds for SD1.1 and SD1.5. We hypothesize that the difference in performance of SD2.1 and SD1.5 is due to the difference in their text encoders (note: differences in performance between SD1.5 and SD2.1 were also reported by several users on online forums (sd2-is-worse-in-face)). Finally, the thresholds for all four datasets are slightly higher for LDM model compared to all the SD models. We hypothesize that this might happen due to the smaller training data size of 400M pairs and leave further investigation for future work.
@@
We also present the plots of the imitation scores as a function of the image frequencies of the concepts for the four datasets. Figures˜5(a) and 5(b) show the imitation graphs of celebrities and classical art styles, respectively for SD1.1. In Figure˜5(a), we observe that the imitation scores for individuals with low image frequencies are close to 0 (left side), and increase as the image frequencies increase (towards the right side). The highest similarity occurs for individuals in the rightmost region of the plot. We also observe a low variance in the imitation scores across prompts for the same concept (the mean of the variance is $0.0003$ with a standard deviation of $0.0005$ across concepts), and note that this variance is independent of the concept frequencies – indicating that the performance of the face embedding model does not depend on the popularity of the individual. Similarly, in Figure˜5(b), we observe that imitation scores for art styles with low image frequencies are low (close to 0.2 on the left side), and increase as the image frequencies increase (towards the right side). The highest similarity occurs for the artists in the rightmost region of the plot. We also notice a low variance across the generation prompts, and the variance does not depend on the popularity of the artist (the mean of the variance is $0.003$ with a standard deviation of $0.003$). We showcase the imitation graphs for all other models in Appendix˜K, which follow similar trends.
(a) Human Face Imitation. The imitation threshold is detected at 399 $\pm$ 85 faces.
(b) Art Style Imitation. The imitation threshold is detected at 304 $\pm$ 56 images.
Figure 5: Human Face and Art Style imitation graphs for SD1.1 using the Celebrities and Classical art style sets. The x-axis is the image frequencies and the y-axis is the imitation score averaged over the five prompts. Concepts with zero image frequencies are shaded in light gray. We show the mean imitation score and its variance over the five image generation prompts for each concept. The red vertical line indicates the imitation threshold found by the change detection algorithm, and the horizontal green line represents the average imitation scores before and after the threshold.
@@
Human Perception Evaluation To determine if the automatic measure of similarity between generated and training images matches human perception, we conduct experiments with human subjects. We ask participants to rate generated images on the Likert scale (likert1932) of 1-5 based on their similarity to real images of celebrities. To avoid any confirmation bias, the participants were not informed of the research objective of this work.
@@
For human face imitation, we conduct this study with 15 participants who were asked to rate 100 (randomly selected) generated images for a set of 40 celebrities on the scale of 1-5. To determine the accuracy of the imitation threshold estimated by MIMETIC2, we randomly select the celebrities such that half of them have image frequencies below the threshold and the other half above it We measure the Spearman’s correlation (zar2005spearman) between the imitation scores computed by the model and the ratings provided by the participants. Due to the variance in perception, we normalize the ratings for each participants. The Spearman correlation between the similarity scores provided by participants and the imitation scores is 0.85, signifying a high quality imitation estimator. We also measure the agreement between the imitation threshold that MIMETIC2 estimates and the threshold that humans perceive. For this purpose, we convert the human ratings to binary values and treat it as the ground truth (any rating of 3 or higher is treated as 1 and ratings less than 3 are treated as 0, indicating successful and unsuccessful imitation, respectively). MIMETIC2 predictions is also binarized, any celebrity with frequency higher than the imitation threshold is treated as 1, otherwise 0. To measure the agreement, we compute the element-wise dot product between these two sets. We find the agreement to be 82.5%, signifying a high degree of agreement for MIMETIC2’s automatically computed threshold.
@@
Evaluation of the Imitation Thresholds Evaluating the accuracy of the real imitation thresholds is expensive, as it requires conducting the counterfactual experiment described in Section˜3 (training $\mathcal{O}(\log m)$ models with different number of images of each concept). To make the evaluation tractable, we finetune a pretrained text-to-image model with concepts that the model has not seen during pretraining and find the number of images required for the model to imitate them. We find that all the concepts require about 250 images for imitation when the models are finetuned on them (further details in Appendix˜H), which is very close to the imitation threshold MIMETIC2 finds for this pretrained model without finetuning (234 images). Based on these results, we conclude that the imitation thresholds estimated by MIMETIC2 are accurate.
Discussion
@@
Overall, we observe that the imitation thresholds are similar across the different image generation prompts, but are domain and model dependent. Importantly, the thresholds computed by MIMETIC2 have a high degree of agreement with human subjects and are supported by the finetuning experiments.
We find that celebrities have a higher imitation threshold than politicians for all models. We hypothesize this happens due to inherent differences in the data distribution between these two sets, which makes it harder to learn the concept of celebrities than politicians. To test this hypothesis, we compute the average number of training images that has a single person for concepts with less than 1,000 images. We find that politicians have about twice the number of single person images compared to celebrities. As such, images that have a single person makes imitation easier for the model (the model can easily associate the face with the name in the caption), thus lowering the imitation threshold for politicians.
@@
We also note the presence of several outliers in both plots, which can be categorized into two types: (1) concepts whose image frequencies are lower than the imitation threshold, but their imitation scores are considerably high; and (2) concepts whose image frequencies are higher than the imitation threshold, but their imitation scores are low. From a privacy perspective, the first kind of outliers are more concerning than the second ones, since the imitation threshold should act as a privacy threshold. When used as a privacy guideline, it is less worrisome if a concept with a concept frequency higher than the threshold is not imitated by the model (false positive), but it would be a privacy violation if a model can imitate a concept with frequency lower than the threshold (false negative). Upon further analysis, we find that for all of the false negative outliers (with the privacy concern), the concepts’ true frequencies are much higher than our estimates. This happens primarily due to naming aliases (a person known by multiple names) that MIMETIC2 does not account for – thereby alleviating the privacy violation concerns (see §7 for a detailed analysis of such outliers).
@@
Finally, we notice that the range of the imitation scores of different domains have different y-axis scales. This is due to the difference in embedding models used in both cases. The face embedding model can distinguish between two faces much better than the art style model can distinguish between two styles (see Appendices˜F and G), and therefore the scores for the concepts on the left side of the imitation threshold is around 0 for face imitation and 0.2 for style imitation. However, the absolute values on the y-axis do not matter for estimating the imitation threshold as long as the trend is similar, which is the case for both domains.